5. A class in statistics contains 10 students, 3 of whom are 19, 4 are 20, 1




is 21, 1 is 24, and 1 is 26. Let X be the average age of the 2 randomly




selected students and derive the probability function for X.






6. A man has four keys in his pocket and, since it is dark, cannot see which




is his door key. He will try each key in turn until he finds the right one.




Let X be the number of keys tried (including the right one) to open the




door. What is the probability function for X?






7. Suppose a fair die is tossed two times. Let X be the larger of the two




faces that appear. Find px(k).






8. Suppose a particle moves along the x-axis beginning at 0. It moves one




integer step to the left or right with equal probability. What is the probability




function of its position after four steps?






9. Five cards are dealt from a standard 52-card deck. Let Y be the number




of red cards that are dealt. What is the probability function for Y ?





1
Expert's answer
2022-02-10T04:03:12-0500

5. We can choose "\dbinom{10}{2}=45" pairs


"(19, 19), (19, 19), (19, 20), (19, 20), (19, 20),"

"(19, 20),(19, 21), (19, 24), (19, 26), (19, 19),"

"(19, 20), (19, 20),(19, 20), (19, 20),(19, 21),"

"(19, 24),(19, 26), (19, 20),(19, 20),(19, 20),"

"(19, 20),(19, 21), (19, 24),(19, 26),(20, 20),"

"(20, 20), (20, 20), (20, 21), (20, 24), (20,26),"

"(20, 20), (20, 20), (20, 21), (20, 24), (20, 26),"

"(20, 20), (20, 21), (20, 24), (20, 26), (20, 21),"

"(20, 24), (20, 26), (21, 24), (21, 26), (24, 26)"

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c}
 & x_i & f_i & f(x_i) \\ \hline
 & 19 & 3& 1/15 \\
 \hdashline
 & 19.5 & 12 & 4/15 \\
 \hdashline
 & 20 & 9 & 3/15 \\
 \hdashline
 & 20.5 & 4 & 4/45 \\
 \hdashline
 & 21.5 & 3 & 1/15 \\
 \hdashline
 & 22 & 4 & 4/45 \\
 \hdashline
 & 22.5 & 4 & 4/45 \\
 \hdashline
 & 23 & 4 & 4/45 \\
 \hdashline
 & 23.5 & 1 & 1/45 \\
 \hdashline
 & 25 & 1 & 1/45 \\
 \hdashline
\end{array}"


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & p(x) \\ \hline
 19 & 1/15 \\
 \hdashline
 19.5 & 4/15 \\
 \hdashline
 20 & 3/15 \\
 \hdashline
 20.5 & 4/45 \\
 \hdashline
 21.5 & 1/15 \\
 \hdashline
 22 & 1/15 \\
 \hdashline
 22.5 & 4/45 \\
 \hdashline
 23 & 4/45 \\
 \hdashline
 23.5 & 1/45 \\
 \hdashline
 25 & 1/45 \\
 \hdashline
\end{array}"


6.

"P(X=1)=\dfrac{1}{4}"

"P(X=2)=(1-\dfrac{1}{4})\dfrac{1}{4-1}=\dfrac{1}{4}"

"P(X=3)=(1-\dfrac{1}{4})(1-\dfrac{1}{4-1})\dfrac{1}{4-2}=\dfrac{1}{4}"

"P(X=4)=(1-\dfrac{1}{4})(1-\dfrac{1}{4-1})(1-\dfrac{1}{4-2})(1)=\dfrac{1}{4}"

"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & 1 & 2 & 3 & 4 \\ \hline
 p(x)& 1/4 & 1/4 & 1/4 & 1/4 \\

 
\end{array}"




7. We have "6^2=36" pairs


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c:c}
 & 1 & 2 & 3 & 4 & 5 & 6 \\ \hline
 1 & 1 & 2 & 3 & 4 & 5 & 6 \\
 \hdashline
 2 & 2 & 2 & 3 & 4 & 5 & 6 \\
 \hdashline
 3 & 3 & 3 & 3 & 4 & 5 & 6 \\
 \hdashline
 4 & 4 & 4 & 4 & 4 & 5 & 6 \\
 \hdashline
 5 & 5 & 5 & 5 & 5 & 5 & 6 \\
 \hdashline
 6 & 6 & 6 & 6 & 6 & 6 & 6 \\
 \hdashline
\end{array}"


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c}
 & x_i & f_i & f(x_i) \\ \hline
 & 1 & 1 & 1/36 \\
 \hdashline
 & 2 & 3 & 1/12 \\
 \hdashline
 & 3 & 5 & 5/36 \\
 \hdashline
 & 4 & 7 & 7/36 \\
 \hdashline
 & 5 & 9 & 1/4 \\
 \hdashline
 & 6 & 11 & 11/36 \\
 \hdashline

\end{array}"


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & p(x) \\ \hline
 1 & 1/36 \\
 \hdashline
 2& 1/12 \\
 \hdashline
 3 & 5/36 \\
 \hdashline
 4 & 7/36 \\
 \hdashline
 5 & 1/4 \\
 \hdashline
 6& 11/36 \\
 \hdashline
\end{array}"


8.


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & p(x) \\ \hline
 -4 & 1/16 \\
 \hdashline
 -2& 1/4\\
 \hdashline
 0 & 3/8 \\
 \hdashline
 +2 & 1/4 \\
 \hdashline
 +4& 1/16\\
 
\end{array}"

9. There are 13 red cards in a deck


"\dbinom{52}{5}=2598960"


"P(X=0)=\dfrac{\dbinom{13}{0}\dbinom{52-13}{5-0}}{\dbinom{52}{5}}=\dfrac{1(575757)}{2598960}"

"=0.2215336"


"P(X=1)=\dfrac{\dbinom{13}{1}\dbinom{52-13}{5-1}}{\dbinom{52}{5}}=\dfrac{13(82251)}{2598960}"

"=0.4114196"


"P(X=2)=\dfrac{\dbinom{13}{2}\dbinom{52-13}{5-2}}{\dbinom{52}{5}}=\dfrac{78(9139)}{2598960}"

"=0.2742797"


"P(X=3)=\dfrac{\dbinom{13}{3}\dbinom{52-13}{5-3}}{\dbinom{52}{5}}=\dfrac{286(741)}{2598960}"

"=0.0815426"


"P(X=4)=\dfrac{\dbinom{13}{4}\dbinom{52-13}{5-4}}{\dbinom{52}{5}}=\dfrac{715(39)}{2598960}"

"=0.0107293"


"P(X=5)=\dfrac{\dbinom{13}{5}\dbinom{52-13}{5-5}}{\dbinom{52}{5}}=\dfrac{1287(1)}{2598960}"

"=0.0004952"


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & p(x) \\ \hline
 1 & 0.4114196 \\
 \hdashline
 2& 0.2742797 \\
 \hdashline
 3 & 0.0815426 \\
 \hdashline
 4 & 0.0107293 \\
 \hdashline
 5 & 0.0004952 \\

\end{array}"